We ran the beam at 1,2 uA on C target and 1,2,3 uA on optics targets (only a few minutes each). I didn't see any noticeable change on the anode currents from the GUI, but we'll analyze epics data more carefully later. A runlist is below, screenshots have been haloged.
6941: 1uA on C 6942: 2uA on C 6943: 1uA on Optics 6943: 2uA on Optics 6945: 3uA on Optics
Only blocks 0 to 191 were on, at 750V. Calorimeter is at 1.5m from the target at 22.3 deg.